OpenStudy (anonymous):

ken leans a 12-foot ladder against his house. He places the ladder so that the base is 5 feet from the house. How far up does the ladder reach?

OpenStudy (anonymous):



OpenStudy (ivettef365):

you can use Pythagorean theorem

OpenStudy (anonymous):

so it would be a^2+b^2=c^2

OpenStudy (ivettef365):

correct

OpenStudy (ivettef365):

so you have 5^2 + h^2 = 12^2 solve for h

OpenStudy (anonymous):

so would i get 25=144

OpenStudy (anonymous):

wait i see what i did wrong it would be h^2+25=144 and you subtract 25 from each side and squareroot 119

OpenStudy (ivettef365):

so you have h^2 = 119 so the answer for h = sqrt119

OpenStudy (anonymous):

yeah so i wouldnt go any further from that

OpenStudy (ivettef365):

or you can put it in decimals which = 10.9